# Describing Globally Distributed Software Architectures for Tax Compliance

### 01 Dec 2023 v1

@@
Abstract
The company-internal reuse of software components owned by organizational units in different countries is taxable. To comply with international taxation standards, multinational enterprises need to consider a geographical perspective on their software architecture. However, there is no viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architecture.
@@ Abstract
In this article, we introduce the reader to the concerns of tax authorities as stakeholders and we investigate how software companies can describe their globally distributed software architectures to tax authorities.
@@
In an in-virtuo experiment, we (1) develop a viewpoint that frames the concerns of tax authorities, (2) create a view of a large-scale, globally distributed microservice architecture from a multinational enterprise, and (3) evaluate the resulting software architecture description with a panel of four tax experts.
@@
The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. However, the resulting view falls short: Although the architecture description reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, taxable, unclear jurisdictions of owners and a potentially insufficient definition of ownership introduce significant noise to the view that limits its usefulness and explanatory power of our software architecture description.
Although our software architecture description provides a solid foundation for the subsequent step, namely valuing software components, we stumbled over several theoretical and practical problems when identifying and defining ownership in distributed teams, which requires further interdisciplinary research.
Index Terms:
architecture description, architecture viewpoint, architecture view, software component, microservice, taxation
I Introduction
@@
Code is not tied to a physical, geographical location due to its intangible nature. The code owner of a software component, however, is. Code ownership refers to the concept of assigning responsibility and accountability for a particular piece of code or software component to an individual, a team, or an organizational unit within an organization.
@@
The geographical location of those software component owners has a legal implication that is too often underestimated: Using software components with owners in different countries is taxable as it constitutes an often implicit form of licensing [1]. To comply with the international standard of taxation [2], multinational enterprises must be able to report a geographical perspective on their software architecture. However, there is no architectural viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architectures which could serve as foundation for a software architecture description.
@@
In this article, we investigate the concerns of tax authorities as stakeholders of globally distributed software architectures and seek an answer to the research question
@@
We address this research question in an in-virtuo experiment with a multinational enterprise aiming to report an architecture description of its globally distributed software architecture. Instead of all involved tax authorities, we build a panel of four tax experts that evaluates the usefulness and completeness of the two core elements of the software architecture description, namely the architecture viewpoint and the architecture view. This experimental design enables us to focus on evaluating emerging technical, procedural, and legal issues in practice while reporting software architecture descriptions to tax authorities without any potential legal implications and consequences—which is crucial in this sensitive context.
@@
Throughout this paper, we use the terminology and definitions of ISO/IEC/IEEE 42010:2022 [3]. In particular, we distinguish between architecture description, architecture view, and architecture viewpoint: An architecture description includes one or more architecture views. An architecture view (or simply, view) addresses one or more of the concerns held by the system’s stakeholders. An architecture view expresses the architecture of the system of interest in accordance with an architecture viewpoint (or simply, viewpoint). There are two aspects to a viewpoint: the concerns it frames for stakeholders and the conventions it establishes on views. Viewpoint conventions can include languages, notations, model kinds, design rules, and/or modelling methods, analysis techniques, and other operations on views.
The remainder of the article is structured as follows: After establishing the background and related work in Section II and discussing our study design as an experimental simulation in Section III, we define the viewpoint (Section IV-B) and how we use it to generate a view for the concern of tax authorities (Section IV-C). In Section V, we present and discuss the evaluation results for our novel viewpoint (Section V-A and the view (Section IV-C) in detail. We conclude the paper with a summary of our findings and the implications for research and practice prior to outlining future work. Figure 1 provides an overview of our study.
@@
Figure 1: An overview of the structure of our study as in-virtuo experiment and its two components: (1) creating a software architecture description with a viewpoint derived from the concerns of tax authorities and a view of an empirical software architecture using the viewpoint and (2) evaluating the software architecture description with a panel of tax experts.
II Related work
@@
This article connects taxation with software architecture distributed across multiple countries. Therefore, we provide related work on globally distributed software development, ownership in the context of software engineering and taxation, and the emerging field of taxation in software engineering in the following subsections.
II-A Globally distributed software development
@@
Globally distributed software development, especially in large enterprises, is a common practice. Tough competition and a lack of resources motivated many software companies to look for ways to optimize their costs by establishing their own subsidiaries in lower-cost premises locally and nationally or even in other countries or by teaming up with allies that can help in delivering products more cost-effectively. As a result, distributed working across national boundaries has become widespread [4]. However, cross-border collaborations are far from being challenge-free. They are recognized to be considerably more complex than the most complicated one-site development endeavors due to the significant communication and coordination overhead. The solutions to the problems of global projects have evolved through the years. They could be characterized by a long journey from initial problematic experiences of collaborative work through strictly independent work in distributed sites to the evolved ways of working across borders.
@@
Architectural modularization was one of the most prominent early solutions to the coordination and communication challenges in global software engineering. In 1968, Conway published his observation that software architecture cannot be separated from the structure of the organization that develops the software, known as Conway’s law [5]. Inspired by this observation, architectural modularization has emerged as a way to minimize dependencies, structure and allocate tasks, and encapsulate the work in a single geographic development site as much as possible so that communication, coordination, and synchronization needs are minimized [6, 7, 8]. Although the modularity approach to coordination has proven very useful, it has many limitations [9, 10, 11]. While some researchers advocate that strict modularization of work is the possibly best way to alleviate the challenges of distributed work, others argue that work in isolation has many disadvantages [9], such as redundant work, cheap and dirty architectural decisions, misplaced functionality [12, 10] and integration problems [13, 14]. Software components are never truly independent and even the best designs are never error-free, while changes are never completely predictable [14].
@@
Further, strict code ownership is also found to introduce communication overhead and delays [11]. For these reasons, and due to the increased use of agile methods, there has been an increasing interest in collaboratively developing software products. This includes a shift from modularization based on geographically distributed locations towards componentization through microservices [15] on a team level, as well as a trend towards establishing collective ownership of source code and other documents to allow everyone in the project to work on any model or artifact [11]. Collaborative work includes keeping track of which team owns and does what (e.g., through Scrum of Scrum fora or organizing cross-unit demos), submitting contributions to other’s components and reviewing each other’s code [1, 16], and fostering reuse of artifacts to reduce redundant work.
II-B Ownership in software engineering and taxation
Ownership plays a crucial role in taxation in general. However, as we learned in the course of this study (see Section V), the definition of ownership in the context of software engineering and the definition of ownership in taxation overlap but are not congruent due to the different context and perspective towards ownership.
Ownership in software engineering, on the one hand, refers to the assignment of responsibility and accountability for specific portions of a software to individuals or teams within a development organization. There are different perspectives on ownership and autonomy in software development organizations. Software development organizations usually adopt the weak ownership principle [17]. This principle entails assigning ownership of specific components (or microservices) to individual teams. In an ideal scenario, the designated team assumes responsibility as the primary (or exclusive) contributor to the associated component. Nevertheless, empirical analysis of real-world instances often reveals a departure from this idealized model [18, 19, 20, 21]. In practice, a given component may offer support to multiple business streams, undergo development by distinct teams, and necessitate the involvement of specialized teams to address non-functional requirements, such as security and reliability. Consequently, while a specific team retains responsibility for the component, the extent of its contribution to the component can exhibit considerable variability [22].
Tax professionals, on the other hand, begin from a civil law perspective and understand ownership in terms of the legal or economic owner regarding the right of exploitation [2]. While the identification of the legal or economic owner is relevant for balance sheet allocations, in the context of transfer pricing, the legal owner is considered (only) as the starting point for a functional and risk analysis that aims for a profit allocation commensurate with the value-adding activities performed by the legal entities within the group, the so-called functional ownership [2].
Describing and discussing the intersection as well as the difference of definitions is beyond the scope of this work and requires future interdisciplinary research.
II-C Taxation
@@
Taxation in software industry has been debated for many decades [23]. The problem with taxing the final result of software engineering, the software product or service, for example, has shown to be challenging to tackle and is still subject to ongoing and broad discussion [24]. This debate has also reached software engineering, the construction of the software product; at first in the context of inner source, use of open source software development practices and the establishment of open source-like communities within an organization [25]. To address the taxation challenges in inner source software development, Buchner and Riehle proposed a cost-based approach [26]. However, Treidler et al. showed that a cost-based approach is not applicable since inner source does not reflect the function and risk profile, the starting point for any transfer pricing discussion, for a cost-based approach [27].
@@
In a previous article [1], we identified the use of software components owned by teams or individuals who represent separate geographically distributed subsidiaries of the same enterprise as type of licensing—a key intangibles in software engineering that are taxable. For an in-depth discussion on the concerns of tax authorities on software architecture, we refer the reader to Section IV-A in which we discuss a simplified example of a multinational enterprise developing software and how tax authorities become a stakeholder to a globally distributed software architecture.
III Study design
@@
We have designed our study as an in-virtuo experiment (or experimental simulation [28]). In such a setup, the behaviors of actors that a researcher aims to observe and measure are natural and rooted in their professional experience and expertise, but the setting in which these behaviors occur has been specifically created for the purposes of the study. That is, without the study, the setting would not exist. Thus, the term “simulation” refers in this context to the artificially created setting that aims to recreate a concrete type of setting in which the “experiment” (i.e., observed behavior) takes place [28].
@@
In our study, we simulate how software companies can describe their globally distributed software architectures for tax compliance. The simulation consists of two parts: (1) creating a view using our viewpoint at a multinational enterprise and (2) the evaluation by a panel of tax experts in which we observe and measure the usefulness and completeness of an architecture description for tax compliance.
@@
On the one hand, the behaviors of actors we aim to observe and measure are all natural: (1) the multinational enterprise fulfilling its reporting obligations regarding their globally distributed software architecture to comply with international taxation standards, (2) the tax advisors representing and consulting companies and creating a line of reasoning, and (3) tax auditors from tax authorities scrutinizing the line of reasoning. We instructed all participants to imagine it to be in the context of a tax audit.
@@
On the other hand, the setting is specifically created for the purpose of this study, also as a means to ensure the anonymity. Obviously, the evaluation of our proposed architecture description in a real tax audit (i.e., a natural, real setting) is not possible: Tax audits are highly sensitive and, if not brought to court, confidential. Our setting must therefore provide a safe space for all participants without any legal implications for the company or the tax authorities.
@@
Although this evaluation is a key pillar of our study, our study goes beyond a traditional judgment study as it also describes the practical issues for creating and using a view based on our novel viewpoint in an empirical case. We discuss our approach to the limitations in Section VII.
Last but not least, we would like to emphasize that an in-depth legal discussion and evaluation of our work is beyond our reach.
III-A Data collection
All involved parties want to remain anonymous including our case company used as a source for creating our view. Therefore, we are not able to describe the case and the data collection in detail. However, we may well argue that our case company is exemplary for a multinational enterprise developing software.
In brief, for creating the view, we extracted the software architecture and the team information from an internal system that extracts the component structure of the system from the code repositories on a daily basis.
For the evaluation of the usefulness and completeness of our software architecture description for its intended purpose, we conducted semi-structured interviews with four tax experts. The main goal of the evaluation is to seek generalizability over the responses rather than generalizability to a population of actors [28]. We ensured privacy and confidentiality for all actors (i.e., the involved companies, tax authorities, and tax advisors) by conducting the interviews separately and removing all company-specific information. We also altered the geographical distribution of the subsidiaries to make our company not identifiable in any sense on the one hand but also to give the reader a better impression of globally distributed software architecture on the other hand.
The interview guidelines are publicly available [^1] https://zenodo.org/records/10214942. The group of four tax experts represent the two opposing parties in a tax audit: internal or external tax advisors representing the company and tax auditors representing the tax authorities.
We conducted two pilot interviews to address didactical challenges since the interviews were conducted with tax experts with no software engineering background. A pair of a co-author with a taxation background and a co-author with a software engineering background conducted the interviews with the tax experts virtually between 2023-10-24 and 2023-11-03. Table I provides an overview of our interviewees and their background. We recorded all interview sessions.
TABLE I: An overview of the interviewees and their backgrounds.
@@ A Tax auditor
III-B Data analysis
As our interview guidelines show, large parts of the interviews were dedicated to providing contextual software engineering understanding for the tax experts. Therefore, we translated and paraphrased only the answers from the interviewees. While one researcher of the interview team analyzed the recorded interviews using thematic analysis [29], the second one ensured the quality of the findings and mappings. After the coding, we again reflected on the findings.
IV A software architecture description for tax compliance
In the following subsections, we provide a gentle introduction to international taxation for software architects and specify the set of conventions for the creation, interpretation, and use of an architecture view to frame the concern of tax authorities.
IV-A International taxation standards as concern and tax authorities as stakeholder

### 04 Mar 2024 v2

@@
Abstract.
Background: The company-internal reuse of software components owned by organizational units in different countries is taxable.
@@ Abstract.
Objective: In this article, we introduce the concerns of tax authorities as stakeholders and investigate how software companies can describe their globally distributed software architectures to tax authorities.
@@
Method: In an experimental simulation, we (1) develop a viewpoint that frames the concerns of tax authorities, (2) create a view of a large-scale, globally distributed microservice architecture from a multinational enterprise, and (3) evaluate the resulting software architecture description with a panel of four tax experts.
@@
Results: The panel found our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. The architecture description reveals that almost 70% of all reuse relationships between the 2560 microservices from our case company are cross-border and, therefore, taxable. However, unclear jurisdictions of owners and potentially insufficient definitions of code ownership and software component introduce significant noise to the view that limits the usefulness and explanatory power of our software architecture description.
Conclusion: Although our software architecture description already provides a solid foundation and reveals the importance of tax compliance in software architectures, we stumbled over several fundamental open questions, forming new frontiers in software engineering.
software architecture, taxation, globally distributed software development
copyright: acmcopyrightisbn: journal: tosemccs: Software and its engineering Software organization and properties
“We don’t do this because it is easy but because we thought it would be easy.”
—Anonymous
1. Introduction
@@
Code is not tied to a physical, geographical location due to its intangible nature. The code owner of a software component, however, is. Code ownership refers to assigning responsibility and accountability for a particular piece of code or software component to an individual, a team, or an organizational unit within an organization.
@@
The geographical location of those software component owners has a legal implication that is too often underestimated: Using software components with owners in different countries is taxable as it constitutes an often implicit form of licensing (Dorner et al., 5555). To comply with the international standard of taxation (OECD, 2022), multinational enterprises must be able to report a geographical perspective on their software architecture. However, there is no architectural viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architecture, which could serve as the foundation for a software architecture description.
@@
In this article, we investigate the concerns of tax authorities as stakeholders of globally distributed software architectures and seek an answer to the research question:
@@
We address this research question in an experimental simulation with a multinational enterprise aiming to report an architecture description of its globally distributed software architecture. Instead of all involved tax authorities, we build a panel of four tax experts evaluating the usefulness and completeness of the two core elements of the software architecture description, namely the architecture viewpoint and the architecture view. This experimental design enables us to focus on evaluating emerging technical, procedural, and legal issues in practice while reporting software architecture descriptions to tax authorities without any potential legal implications and consequences—which is crucial in this sensitive context.
@@
Throughout this paper, we use the terminology and definitions of ISO/IEC/IEEE 42010:2022 (ISO/IEC/IEEE42010:2022, 2022). In particular, we distinguish between architecture description, architecture view, and architecture viewpoint: An architecture description includes one or more architecture views. An architecture view (or simply, view) addresses one or more of the concerns held by the system’s stakeholders. An architecture view expresses the architecture of the system of interest in accordance with an architecture viewpoint (or simply, viewpoint). There are two aspects to a viewpoint: the concerns it frames for stakeholders and the conventions it establishes on views. Viewpoint conventions can include languages, notations, model kinds, design rules, and/or modelling methods, analysis techniques, and other operations on views.
The remainder of the article is structured as follows: After establishing the background and related work in Section 2 and discussing our study design as an experimental simulation in Section 3, we define the viewpoint (Section 4.2) and how we use it to generate a view for the concern of tax authorities (Section 4.3). In Section 5, we present and discuss the evaluation results for our novel viewpoint (Section 5.1 and the view (Section 4.3) in detail. We conclude the paper with a summary of our findings and the implications for research and practice before outlining future work. Figure 1 provides an overview of our study.
@@
Figure 1. An overview of the structure of our study as experimental simulation and its two components: (1) creating a software architecture description with a viewpoint derived from the concerns of tax authorities and a view of an empirical software architecture using the viewpoint and (2) evaluating the software architecture description with a panel of tax experts.
2. Related work
@@
This article connects taxation with software architecture distributed across multiple countries. Therefore, we provide related work on globally distributed software development and the emerging field of taxation in software engineering in the following subsections.
2.1. Globally distributed software development
@@
Globally distributed software development, especially in large enterprises, is a common practice. Tough competition and a lack of resources motivated many software companies to look for ways to optimize their costs by establishing their own subsidiaries in lower-cost premises locally and nationally or even in other countries or by teaming up with allies that can help in delivering products more cost-effectively. As a result, distributed working across national boundaries has become widespread (Ruhe and Wohlin, 2014). However, cross-border collaborations are far from being challenge-free. They are recognized to be considerably more complex than the most complicated one-site development endeavors due to the significant communication and coordination overhead. The solutions to the problems of global projects have evolved through the years. They could be characterized by a long journey from initial problematic experiences of collaborative work through strictly independent work in distributed sites to the evolved ways of working across borders.
@@
Architectural modularization was one of the most prominent early solutions to the coordination and communication challenges in global software engineering. In 1968, Conway published his observation that software architecture cannot be separated from the structure of the organization that develops the software, known as Conway’s law (Conway, 1968). Inspired by this observation, architectural modularization has emerged as a way to minimize dependencies, structure and allocate tasks, and encapsulate the work in a single geographic development site as much as possible so that communication, coordination, and synchronization needs are minimized (Cataldo et al., 2008; Herbsleb and Mockus, 2003; Herbsleb and Grinter, 1999a). Although the modularity approach to coordination has proven very useful, it has many limitations (Cataldo and Herbsleb, 2008; Šmite, 2014; Clerc et al., 2007). While some researchers advocate that strict modularization of work is the possibly best way to alleviate the challenges of distributed work, others argue that work in isolation has many disadvantages (Cataldo and Herbsleb, 2008), such as redundant work, cheap and dirty architectural decisions, misplaced functionality (Tureček et al., 2010; Šmite, 2014) and integration problems (Kwan et al., 2011; Herbsleb and Grinter, 1999b). Software components are never truly independent and even the best designs are never error-free, while changes are never completely predictable (Herbsleb and Grinter, 1999b).
@@
Further, strict code ownership is also found to introduce communication overhead and delays (Clerc et al., 2007). For these reasons, and due to the increased use of agile methods, there has been an increasing interest in collaboratively developing software products. This includes a shift from modularization based on geographically distributed locations towards componentization through microservices (Newman, 2021) on a team level, as well as a trend towards establishing collective ownership of source code and other documents to allow everyone in the project to work on any model or artifact (Clerc et al., 2007). Collaborative work includes keeping track of which team owns and does what (e.g., through Scrum of Scrum fora or organizing cross-unit demos), submitting contributions to other’s components and reviewing each other’s code (Dorner et al., 5555; Šmite et al., 2023), and fostering reuse of artifacts to reduce redundant work.
2.2. Taxation in software engineering
@@
Taxation in software industry has been debated for many decades (OECD, 2015). The problem with taxing the final result of software engineering, the software product or service, for example, has shown to be challenging to tackle and is still subject to ongoing and broad discussion (Olbert and Spengel, 2017). This debate has also reached software engineering, the construction of the software product; at first in the context of inner source, use of open source software development practices and the establishment of open source-like communities within an organization (Capraro, 2020). To address the taxation challenges in inner source software development, Buchner and Riehle proposed a cost-based approach (Buchner and Riehle, 2022). However, Treidler et al. showed that a cost-based approach is not applicable since inner source does not reflect the function and risk profile, the starting point for any transfer pricing discussion, for a cost-based approach (Treidler et al., 2024).
@@
Dorner et al. identified the use of software components owned by teams or individuals who represent separate geographically distributed subsidiaries of the same enterprise as type of licensing—a key intangible in software engineering that is taxable (Dorner et al., 5555). For an in-depth discussion on the concerns of tax authorities on software architecture, we refer the reader to Section 4.1 in which we discuss a simplified example of a multinational enterprise developing software and how tax authorities become a stakeholder to a globally distributed software architecture.
3. Study design
@@
We designed our study as an experimental simulation (Stol and Fitzgerald, 2018). In such a setup, the behaviors of actors that a researcher aims to observe and measure are natural and rooted in their professional experience and expertise, but the setting in which these behaviors occur has been specifically created for the purposes of the study. That is, without the study, the setting would not exist. Thus, the term “simulation” refers in this context to the artificially created setting that aims to recreate a concrete type of setting in which the “experiment” (i.e., observed behavior) takes place (Stol and Fitzgerald, 2018).
@@
In our study, we simulate how software companies can describe their globally distributed software architectures for tax compliance. The simulation consists of two parts: (1) creating a view using our viewpoint at a multinational enterprise and (2) evaluating by a panel of tax experts in which we observe and measure the usefulness and completeness of an architecture description for tax compliance.
@@
On the one hand, the behaviors of actors we aim to observe and measure are all natural: (1) the multinational enterprise fulfilling its reporting obligations regarding its globally distributed software architecture to comply with international taxation standards, (2) the tax advisors representing and consulting companies and creating a line of reasoning, and (3) tax auditors from tax authorities scrutinizing the line of reasoning. We instructed all participants to imagine it to be in the context of a tax audit.
@@
On the other hand, the setting is specifically created for the purpose of this study, also as a means to ensure anonymity and immunity. Obviously, evaluating our proposed architecture description in a real tax audit (i.e., a natural, real setting) is not possible: Tax audits are highly sensitive and, if not brought to court, confidential. Our setting must, therefore, provide a safe space for all participants without any legal implications for the company or the tax authorities.
@@
Although this evaluation is a crucial pillar of our study, our study goes beyond a traditional judgment study as it also describes the practical issues for creating and using a view based on our novel viewpoint in an empirical case. We discuss our approach to the limitations in Section 7.
Last but not least, we would like to emphasize that an in-depth legal and economic discussion and evaluation of our work is beyond our reach. We further discuss this limitation and its implications for research in Section 8.
3.1. Data collection
All involved parties want to remain anonymous, including our case company, which was used as a source for creating our view. Therefore, we are not able to describe the case and the data collection in detail. However, we may well argue that our case company is exemplary for a multinational enterprise developing software.
In brief, to create the view, we extracted the software architecture and the team information from an internal system that extracts the component structure of the system from the code repositories on a daily basis.
To evaluate the usefulness and completeness of our software architecture description for its intended purpose, we conducted semi-structured interviews with four tax experts. The main goal of the evaluation is to seek generalizability over the responses rather than generalizability to a population of actors (Stol and Fitzgerald, 2018). We ensured privacy and confidentiality for all actors (i.e., the involved companies, tax authorities, and tax advisors) by conducting the interviews separately and removing all company-specific information. We also altered the geographical distribution of the subsidiaries to make our company not identifiable in any sense on the one hand but also to give the reader a better impression of globally distributed software architecture on the other hand.
The interview guidelines are publicly available [^1] https://zenodo.org/records/10214942. The group of four tax experts represents the two opposing parties in a tax audit: internal or external tax advisors representing the company and tax auditors representing the tax authorities.
We conducted two pilot interviews to address didactical challenges since the interviews were conducted with tax experts with no software engineering background. A pair of a co-author with a taxation background and a co-author with a software engineering background conducted the interviews with the tax experts virtually between 2023-10-24 and 2023-11-03. Table 1 provides an overview of our interviewees and their backgrounds. We recorded all interview sessions.
@@ A Tax auditor
Table 1. An overview of the interviewees and their backgrounds.
3.2. Data analysis
As our interview guidelines show, large parts of the interviews were dedicated to providing contextual software engineering understanding for the tax experts. Therefore, we translated and paraphrased only the answers from the interviewees. While one researcher of the interview team analyzed the recorded interviews using thematic analysis (Clarke and Braun, 2017), the second one ensured the quality of the findings and mappings. After the coding, we again reflected on the findings.
4. A software architecture description for tax compliance
In the following subsections, we provide a gentle introduction to international taxation for software architects and specify the set of conventions for creating, interpreting, and using an architecture view to frame the concern of tax authorities.
4.1. International taxation standards as concern and tax authorities as stakeholder